i have been away from school this past week. can you help?
Given Z=g(y) = 2y+3 and y=f(x) = x-1 then gof =

I came up with 2x-1 it is a homework question thanks

Answer by stanbon(48558)

g(y) = 2y+3 and y=f(x) = x-1 then gof = ?
----------
gof means g[f(x)]
=g[x-1]
=2(x-1)+3
=2x-2+3
=2x+1
============
Cheers,
Stan H.